OpenStudy (ammarah):

Find the vertex of this quadratic function: r(x)=-4x^2+16x-15

OpenStudy (ammarah):

@satellite73

OpenStudy (matlee):

@Zarkon

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}\)

OpenStudy (anonymous):

in your case that is \[-\frac{16}{2\times (-4)}=2\]

OpenStudy (anonymous):

second coordinate of the vertex is what you get when you replace \(x\) by the first coordinate

OpenStudy (ammarah):

no i dont get how to set up the equation like what i did was add 15 to both sides

OpenStudy (ammarah):

@ganeshie8

OpenStudy (shinalcantara):

\[r(x) =-4x^2 + 16x - 15\] ------------------------ Use vertex form \[y = a(x-h)^2 + k\] where (h,k) are the coordinates of the vertex ------------------------ Group thru parenthesis \[r(x) = (-4x^2 + 16x) - 15\] ------------------------ factor out -4 \[r(x) = -4(x^2 + 4x) - 15\] ------------------------ Complete the square by adding and subtracting 16 \[r(x) = -4(x^2 + 4x + 4) - 15 + 16\] Notice that we only add 4 in the parenthesis since when -4 is distributed, it will become -16 ------------------------- \[r(x) = -4(x^2 + 4x + 4) - 15 + 16\] \[r(x) = -4(x+2) + 1\] h=-2, k=1 ------------------------- Therefore the coordinates of the vertex is (-2,1).

OpenStudy (shinalcantara):

* \[r(x)=-4(x+2)^2+1\]